OpenStudy (anonymous):

Find the horizontal asymptote of f(x)=(x^2+8x-2)/(x-2)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}(x^2+8x-2)/(x-2)= \infty\]

OpenStudy (anonymous):

no horizontal asymptote

OpenStudy (anonymous):

thanx